State De Moivre’s Theorem. (2)
9.2 Express cos 5θ and sin 4θ as polynomials in terms of sin θ and cos θ. (8)
9.3 Let w be a negative real number, z a 6
th root of w.
(a) Show that z (k) = ρ
1
6
-

cos
1
Expert's answer
2020-04-27T18:18:23-0400

9.2 As per the given question,

As per demoiver's theorem,

"(\cos\theta+i\sin \theta)^5=\cos 5\theta +i\sin5\theta"

Now, let the binomial expansion of the left part,

"\cos^5\theta+5\cos^4\theta(i\sin\theta)+10\cos^3\theta(i\sin\theta)^2+10\cos^2\theta(i\sin\theta)^3+5\cos\theta(i\sin\theta)^4+(i\sin\theta)^5"

Now, equating the real and imaginary part of the above,

"\\cos 5\\theta =\\cos^5\\theta \u2212 10\u2009\\cos^3\u2009\\theta\u2009\\sin2\\theta + 5\u2009\\cos\\theta\u2009\\sin4\\theta"

Similarly,

"(\cos\theta+i\sin \theta)^4=\cos 4\theta +i\sin4\theta"

Now the expansion lhs of the equation,

"\cos^4\theta+4\cos^3\theta(i\sin\theta)+6\cos^2\theta(i\sin\theta)^2+4\cos\theta(i\sin\theta)^3+(i\sin\theta)^4"

Now, comparing the real and imaginary part,

"\sin 4\theta=4\cos^3\theta \sin \theta +4 \cos\theta \sin^3\theta"


9.3 As per the question, data is insufficient to answer 9.3, I am taking some assumptions,

that it is z is the complex number,

the polar form of the complex number, z= a("\cos\theta+i\sin \theta" )

As per the question, it is given that w is the negative real number so,"(w)=(2k+1)\pi"

k=6,

"w=13\pi"

"z^6=a^6(\cos\theta+i\sin\theta)^6"

As per the demovier's rule

"=a^6(\cos6\theta+i\sin6\theta)"

6th real root ="\cos 6\theta=\cos 13\pi" =-1


"z=re^{i\u03b8}" and convert1to polar form to get

"z^6=w"

z and w in polar form "z=re^{i\u03b8},w=se^{i\u03c6}" . Then "z^6=w"

becomes:"(re^{i\u03b8})^6=r^6e^{in\u03b8}=se^{i\u03c6}"

"r^6=s"

"e^{6\theta}=e^{i\phi}"

"r^6=s"

"r=s^{1\/6}"

"θ=φ/n+2πl/n=2\pi/6=\pi/3+2πl/6"

6th root of the "=re^{i(\pi/3+2πl/6)}"

As data is not insufficient and incomplete so have chooses some approximation.